Find all real or imaginary solutions to the equation.
w^2/3 = 4

Answer by stanbon(75887)

Find all real or imaginary solutions to the equation.
w^2/3 = 4
(w^(2/3))^(3/2) = 4^(3/2)
w = 2^3
w = 8
============
Cheers,
Stan H.